OpenStudy (anonymous):

Part A: Max rented a motorbike at $465 for 5 days. If he rents the same motorbike for a week, he has to pay a total rent of $625. Write an equation in the standard form to represent the total rent (y) that Max has to pay for renting the motorbike for x days. (4 points) Part B: Write the equation obtained in Part A using function notation. (2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (anonymous):

what i already have part A:625-465=$160 he has to pay for 2 days=$160 rent paid for 1 day=260/2=$80 amount paid for 5days=8*5=$400 But he paid $465 Difference 465-400=$65 so every one has to pay $65+rent for the no. of days total rent for x days y=65+80 x (dollars) part B: f(x)=65+80 x part C: if thats right i just need help with c.. please

OpenStudy (anonymous):

y = amount x = days min amount = 65+80 = 145 so x = 1, 2, 3, 4, 5 and so on. 1 unit = 1 day and y = 145, 225, 305 and so on. ! unit = $80

OpenStudy (anonymous):

can you help

OpenStudy (anonymous):

is that c ?

OpenStudy (anonymous):

thats for c part.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

anytime :)